Nash Enterprises purchased a delivery truck on January 1, 2017, at a cost of $28,080. The truck has a useful life of 7 years with an estimated salvage value of $6,240. The straight-line method is used for book purposes. For tax purposes, the truck, having an MACRS class life of 7 years, is classified as 5-year property; the optional MACRS tax rate tables are used to compute depreciation. In addition, assume that for 2017 and 2018 the company has revenues of $208,000 and operating expenses (excluding depreciation) of $135,200.

MACRS Depreciation Rates by Class of Property

Recovery
Year

3-year
  (200% DB)

5-year
(200% DB)

7-year
(200% DB)

10-year
(200% DB)

15-year
(150% DB)

20-year
(150% DB)

1        

33.33      20.00      14.29      10.00      5.00      3.750     

2        

44.45      32.00      24.29      18.00      9.50      7.219     

3        

14.81*    19.20      17.49      14.40      8.55      6.677     

4        

      7.41      11.52*    12.49      11.52      7.70      6.177     

5        

11.52      8.93*    9.22      6.93      5.713     

6        

5.76      8.92      7.37      6.23      5.285     

7        

8.93      6.55*     5.90*    4.888     

8        

4.46      6.55      5.90      4.522     

9        

6.56      5.91      4.462*   

10        

6.55      5.90      4.461     

11        

3.28      5.91      4.462     

12        

5.90      4.461     

13        

5.91      4.462     

14        

5.90      4.461     

15        

5.91      4.462     

16        

2.95      4.461     

17        

4.462     

18        

4.461     

19        

4.462     

20        

4.461     

21        

2.231     

*Switchover to straight-line depreciation.

Compute income before income taxes for 2017 and 2018.

2017

2018

Income before income taxes

Compute taxable income for 2017 and 2018. (Round answers to 0 decimal places, e.g. 45,892.)

Taxable income for 2017

  

Taxable income for 2018

Determine the total depreciation to be taken over the useful life of the delivery truck for both book and tax purposes.

Book purposes   
Tax purposes
Answer #1

computation of income before income tax

particular 2017 2018

revenues 208000 208000

less: operating expense 135200 135200

less: depreciation 3120 3120

income before income tax 69680 69680

computation of taxable income

particulars 2017 2018

revenues 208000 208000

less: operating expenses 135200 135200

less: depreciation for tax purpose 5616 7188

taxable income 67184 65612

computation of total depreciation over useful life of the truck

for book purpose = (28080-6240)/7= 3120 for 1 year

total depreciation = 3120*7= 21840

for tax purpose

for year 1 28080* 20% = 5616

year 2 22464*32% = 7188

year 3 15276*19.20%= 2933

year 4 12343*11.52%= 1422

year 5 10921*11.52%=1258

year 6 9663*5.76%= 556

total depreciation= 18973